OpenStudy (anonymous):

a triangle has side lengths of 6 8 and 9 what type of triangle is it a.obtuse b.right c.acute d.equiangular?

OpenStudy (anonymous):

an equiangular triangle has equal sides (equilateral). So you can rule that out. To figure out the answer, you can take the longest side and pretend it's the hypotenuse in a right triangle and try out the Pythagorean formula. If a^2 + b^2 = c^2, it's a right triangle. If a^2 + b^2 > c^2, it's acute. If a^2 + b^2 < c^2, it's obtuse.

OpenStudy (anonymous):

c) acute

OpenStudy (anonymous):

an equiangular triangle is an equilateral triangle and has all sides of equal length-- so that is ruled out 6^2 + 8^2 = 36 + 64 = 100 and 9^2 is 81 so sum of squares of two sides is not equal to the square of the third side, hence not right angled triangle.... and as hickninja said for the obtuse angle triangle....

OpenStudy (anonymous):

so its an acute triangle?

OpenStudy (anonymous):

if a^2 + b^2 > c^2, it's acute. a^2 + b^2 = 100. c^2 is 81.

OpenStudy (anonymous):

yes i said that earlier too... its acute angled triangle

OpenStudy (anonymous):

It's a scalene triangle with all interior angles acute. If the longest side were 10, you would have a right triangle. The longest side is shorter than 10, so the angle opposite it is acute. Since it is the largest of the interior angles, they are all acute.